OpenStudy (anonymous):

can someone explain how to find the answer please? What is the axis of symmetry for f(x) = −3x2 + 18x − 7? x = −3 x = 3 x = −6 x = 6

OpenStudy (anonymous):

find the vertex of \(f(x) = −3x^2 + 18x − 7 \) or just compute \(-\frac{b}{2a}\) with \(a=-3,b=18\)

OpenStudy (anonymous):

axis of symmetry for \[y=ax^2+bx+c\] is \[x=-\frac{b}{2a}\]

OpenStudy (anonymous):

so all i do is plug in the points, (-3,18), and solve?

OpenStudy (anonymous):

not sure what you mean, it is not a point, it is a number

OpenStudy (anonymous):

i.e. \[-\frac{b}{2a}=-\frac{18}{2\times (-3)}\] is what you need to compute for the axis of symmetry

OpenStudy (anonymous):

the axis of symmetry is a vertical line the equation for a vertical line looks like \(x=number\) in your case the number is \(3\) so the axis of symmetry is \(x=3\)

OpenStudy (anonymous):

okay thank you so much!

OpenStudy (anonymous):

yw